# WiSegRT: Dataset for Site-Specific Indoor Radio Propagation Modeling with 3D Segmentation and Differentiable Ray-Tracing

### 18 Dec 2023 v1

@@ V-B Indoor Radio Sensing
As EM wave has been deployed in radars for sensing from 1930s, employing radio for indoor sensing is a natural progression due to the ubiquity of wireless devices. Numerous innovative studies have leveraged radio waves for indoor sensing applications, like monitoring human pose and even vital signs. As we provide the 3D model of the environment, it support further scenario customization, including the addition of human figures, editing their actions, and even varying levels of perspiration as it largely change the EM properties of skin and clothing. By understanding how radio interact with target and subsequently reach the Rx, we can build more physics-informed approaches rather than just end-to-end models from the noisy measurements.
@@ VI Conclusions
In this paper, we introduced the WiSegRT, a precise indoor radio propagation dataset. Compared with the simple settings deployed by previous works to generate dataset, the WiSegRT dataset provides an enhanced representation of the complex characteristics of real-world radio propagation. With the virtual-reality-like scene models, WiSegRT supports various ML-based wireless tasks, such as the vision-based radio tracing and ray-level indoor radio-sensing discussed in the application section. Future works will focus on developing model editing libraries to support digital-twin-styled tasks.

### 31 Dec 2023 v2

@@ V-B Indoor Radio Sensing
As EM wave has been deployed in radars for sensing from 1930s, employing radio for indoor sensing is a natural progression, driven by the widespread adoption of wireless devices. Many pioneering studies have leveraged radio waves for indoor sensing applications, such as monitoring human pose and even vital signs. As we provide the 3D model of the environment, it supports further scenario customization, including the addition of human figures, editing their actions, and even varying levels of perspiration as it largely changes the EM properties of skin and clothing. By understanding how radio waves interact with target and subsequently reach the Rx, we can build more physics-informed approaches rather than relying solely on end-to-end models from the noisy measurements.
@@ VI Conclusions
In this paper, we introduced the WiSegRT, a precise indoor radio propagation dataset. Compared with the simple settings deployed by previous works to generate dataset, the WiSegRT dataset provides an enhanced representation of the complex characteristics of real-world radio propagation. With the virtual-reality-like scene models, WiSegRT supports various ML-based wireless tasks, such as the vision-based radio tracing and ray-level indoor radio-sensing. Future works will focus on developing model editing libraries to support digital-twin-styled tasks.